OpenStudy (anonymous):

Hena stores coffee powder in a steel container with a capacity of 82.5 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 4.2 inches, what is its radius?

OpenStudy (anonymous):

volume of a cylinder = Base Area x Height Base Area= area of circle \[V = \pi r^2 h\] congruent will mean they have equal volume let V = 82.5 and let h = 4.2 \[82.5= \pi \times 4.2 \times r^2\]

OpenStudy (anonymous):

SO whats the answer? :\

OpenStudy (anonymous):

82.5/π/4.2=6.25 squared =2.5